You purchased 5,600 shares in the New Pacific Growth Fund on January 2, 2016, at an offering price of $45.90 per share. The front-end load for this fund is 4 percent, and the back-end load for redemptions within one year is 2 percent. The underlying assets in this mutual fund appreciate (including reinvested dividends) by 5 percent during 2016, and you sell back your shares at the end of the year. If the operating expense ratio for the New Pacific Growth Fund is 1.75 percent, what is your total return from this investment? (Assume that the operating expense is netted against the fund’s return.) (A negative value should be indicated by a minus sign. Do not round intermediate calculations. Enter your answer as a percent rounded to 2 decimal places.)
Answer #1

Total return (amount) = net sale proceeds - initial investment

Initial investment

Initial investment is $45.90 per share.

Number of shares purchased is 5600.

Total initial investment = $45.90 * 5600 = $257,040.

Net sale proceeds

Net initial investment = gross investment * (1 - front-end load)

Net initial investment = $257,040 * (1 - 4%) = $246,758.40.

Ending value of assets = Net initial investment * (1 + net rate of return).

net rate of return = gross rate of return - expense ratio.

net rate of return = 5% - 1.75% = 3.25%.

Ending value of assets =  $246,758.40 * (1 + 3.25%) = $254,778.05.

Net sale proceeds = ending value of assets * (1 - backend load)

Net sale proceeds = $254,778.05 * (1 - 2%) = $249,682.49.

Total return = net sale proceeds - initial investment

Total return = $249,682.49 - $257,040  

Total return (amount) = -$7,357.51.

Total return (%) = Total return (amount) / initial investment

Total return (%) = -$7,357.51 / $257,040  

Total return (%) = -2.86%